OpenStudy (anonymous):

Find the quadratic equation that fits the following set of data points. {(-1,4),(3,-36),(-5,12)} Write the equation in standard form ax^2+bx+c y=

OpenStudy (anonymous):

three variables, three equations replace x by -1, set the result equal to 4 \[a-b+c=4\] replace x by 3 set result equal to -36 \[9a+2b+c=36\] replace x by -5, set result equal to 12 \[25a-5b+c=12\]

OpenStudy (anonymous):

solve the system \[a-b+c=4\] \[9a+2b+c=-36\] \[25a-5b+c=12\]

OpenStudy (anonymous):

typo there, second equation should be \[9a+3b+c=-36\]

OpenStudy (anonymous):

that's better. \(a=-1,b=-8,c=-3\)